A string vibrate at a frequency of 250hz if the distance between the 1 first and the 5th antinode is 40cm. Calculate the velocity of vibration??
1
Expert's answer
2018-06-25T11:17:08-0400
First, 40 cm – is equal to 0.4 m. We know, that distance between antinodes is a half of wavelength.
So:
(5-1)×λ/2=0.4
4×λ/2=0.4
2λ=0.4
λ=0.2
We also know that velocity of vibration is v=λ, where λ – is a wavelength, and  - is frequency.
So:
v=0.2×250=50 m/s